In a sample of 150 ilocanos, 65% wished that they were rich. In a sample of 260 Visayas, 60% wished that they were rich. At š›¼ = 0.05, is there a difference in the proportions? Find the 95% confidence interval for the difference of two proportions


1
Expert's answer
2021-12-01T14:21:11-0500

a) The value of the pooled proportion is computed as


"\bar{p}=\dfrac{X_1+X_2}{n_1+n_2}=\dfrac{0.65(150)+0.6(260)}{150+260}=0.6183"

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p_1=p_2"

"H_1:p_1\not=p_2"

This corresponds to a two-tailed test, and a z-test for two population proportions will be used.

Based on the information provided, the significance level isĀ "\alpha = 0.05," and the critical value for a two-tailed test isĀ "z_c = 1.96."

The rejection region for this two-tailed test isĀ "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}_1-\hat{p}_2}{\sqrt{\bar{p}(1-\bar{p})(1/n_1+1/n_2)}}"

"=\dfrac{0.65-0.6}{\sqrt{0.6183(1-0.6183)(1/150+1/260)}}"

"=1.0038"

Since it is observed thatĀ "|z| = 1.0038 \le1.96= z_c," it is then concluded thatĀ the null hypothesis is not rejected.

Using the P-value approach:

The p-value isĀ "p =2P(Z>1.0038)= 0.315475," and sinceĀ "p= 0.315475>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportionĀ "p_1"Ā is different thanĀ "p_2," at theĀ "\alpha = 0.05" significance level.


b) The critical value forĀ "\alpha = 0.05"Ā isĀ "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\hat{p}_1-\hat{p}_2-z_c\sqrt{\dfrac{\hat{p}_1(1-\hat{p}_1)}{n_1}+\dfrac{\hat{p}_2(1-\hat{p}_2)}{n_2}},"

"\hat{p}_1-\hat{p}_2+z_c\sqrt{\dfrac{\hat{p}_1(1-\hat{p}_1)}{n_1}+\dfrac{\hat{p}_2(1-\hat{p}_2)}{n_2}})"

"=(0.65-0.6-1.96\sqrt{\dfrac{0.65(1-0.65)}{150}+\dfrac{0.6(1-0.6)}{260}},"

"0.65-0.6+1.96\sqrt{\dfrac{0.65(1-0.65)}{150}+\dfrac{0.6(1-0.6)}{260}})"

"\approx(-0.047, 0.147)"

Therefore, based on the data provided, theĀ 95 %Ā confidence interval for the difference between the population proportionsĀ "p_1 - p_2" isĀ "-0.047 < p_1 - p_2 < 0.147," which indicates that we areĀ 95%Ā confident that the true difference between population proportions is contained by the intervalĀ "(-0.047, 0.147)."